# Work extraction with feedback control using limited resources

### v1

@@ V.2 Work extraction from random initial distributions
A second related but different scenario of interest is the case in which the initial distribution $\rho_{0}(\mathbf{x})$ of the system is not completely known to the feedback controller. This includes and generalizes the case of the cheeseburger effect considered in Sec. III.3. Uncertainty in the initial distribution amounts to a random macrostate, from which the randomness of the microstates given the macrostate is subsidiary. In this setting, the measurement is about the distribution (i.e., information regarding which of several possible distributions may describe the system), rather than of the distribution (i.e., a noisy measurement of the system state).
In particular, we assume that the initial distribution $\rho_{0}^{(\alpha)}$ is sampled from a finite set labeled by $\alpha$ according to some prefixed probability $p_{\alpha}$, with $\sum_{\alpha}p_{\alpha}=1$. In this case we consider that the feedback controller performs some (imperfect) measurement of the system concerning the value of $\alpha$ of the original distribution, rather than the state $\mathbf{x}$. Then the conditional probability capturing the effect of external noise is of the form ${P}(m|\alpha)$, where now the result of the measurement $m$ is a single number related to the values of the parameter $\alpha$, occurring with (marginal) probability $p(m)=\sum_{\alpha}P(m|\alpha)p_{\alpha}$. Such a situation is to be compared with the cases in which no information about the parameter $\alpha$ is available, implying an effective mixture $\bar{\rho}_{0}=\sum_{\alpha}p_{\alpha}\rho_{0}^{(\alpha)}$, and to the case in which perfect knowledge of the initial distribution $p_{0}^{(\alpha)}$ is assumed.
After performing the noisy measurement of $\alpha$, the controller can gain knowledge of the initial distribution of the system. The conditional probability of obtaining a possible $\alpha$ value given the measurement result $m$ is ${P}(\alpha|m)={P}(m|\alpha)p_{\alpha}/p(n)$, from which the updated system distribution ${\rho}_{X|m}(x)$ can be computed as:
@@
We now consider the combination of the situations of random initial distributions as described above and control limitations. Since each measurement outcome $m$ is unambiguously related to an optimal post-quench equilibrium distribution $\rho_{X|m}(\mathbf{x})$, our previous logic will apply. If there are $M$ possible measurements $m\in\{1,...,M\}$, and $N$ possible quenching Hamiltonians, $n\in\{1,...,N\}$, where $N<M$, then as before we need to design a repertoire and lookup table $n(m)$ mapping from the measurement results to the associated quenches in $\mathcal{H}$. Given a partition =$\mathcal{L}=\{\ell(n)\}_{n=1}^{N}$, similar calculations to those of Appendix A will apply to obtain an optimal $\rho_{n}^{\rm opt}(\mathbf{x})$ analogous Eq. (28). Then likewise the average entropy production under control limitations would read:
$\langle S_{\mathrm{tot}}\rangle=\sum_{\alpha}p_{\alpha}\sum_{m}P(m|\alpha)D(\rho_{0}^{(\alpha)}||\rho_{n^{*}(m)}^{\rm opt})$ (37)
which may exhibit a representation in terms of JSD akin to Eq (30). The measurement of $\alpha$ can provide macroscopic information of significant energetic value; see Sec. III.3.
@@ VI Conclusions and Discussion
In addition, such limitations are abundant in social systems. For example, many modern hunter-gatherer groups maintain social norms that help facilitate collective behavior, including the deterrence of free-riders and the encouragement of demand sharing (i.e., sharing food according to need; Lewis et al. (2014)). Thus, when an individual hunter-gatherer finds food or is successful in a hunt, their energy extraction actions are limited: they might eat some of the food immediately, but must save the rest to share with their camp. Further, most hunter-gatherer populations do not generally store or preserve food Armit and Finlayson (1992). They consume it soon after collecting it or they loose out on extracting energy from it. In contrast, individuals living in farming societies rely on being able to choose among multiple actions. Namely, to consume food immediately, or store it for later use. The increasingly complex technology available to farming and urbanized societies suggests that having access to a larger action repertoire may be crucial to explaining the massive increase in per-person energy use between the ancient and modern world Krausmann et al. (2016).
@@ Acknowledgements
H. H., D. W., A. S., and C. K. acknowledge support from the Santa Fe Institute. D. W. acknowledges support from US NSF Grant CCF-2221345. G.  M. acknowledges support from the “Ramón y Cajal” program (No. RYC2021-031121-I), the CoQuSy project (No. PID2022-140506NB-C21), and the María de Maeztu project (No. CEX2021-001164-M) funded by MCIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR.

### v2

@@ V.2 Work extraction from random initial distributions
A second related but different scenario of interest is the case in which the initial distribution $\rho_{0}(\mathbf{x})$ of the system is not completely known to the feedback controller.
Uncertainty in the initial distribution amounts to a random macrostate, from which the randomness of the microstates given the macrostate is subsidiary. In this setting, the measurement is about the distribution (i.e., information regarding which of several possible distributions may describe the system), rather than of the distribution (i.e., a noisy measurement of the system microstate).
In particular, we assume that the initial distribution $\rho_{0}^{(\alpha)}(\mathbb{x})$ is sampled from a finite set labeled by $\alpha$ according to some prefixed probability $p_{\alpha}$, with $\sum_{\alpha}p_{\alpha}=1$. In this case we consider that the feedback controller performs some (imperfect) measurement of the system concerning the value of $\alpha$ of the original distribution, rather than the state $\mathbf{x}$. Then the conditional probability capturing the effect of external noise is of the form ${p}(m|\alpha)$, where now the result of the measurement $m$ is a single number related to the values of the parameter $\alpha$, occurring with (marginal) probability $p(m)=\sum_{\alpha}p(m|\alpha)p_{\alpha}$. Such a situation is to be compared with the cases in which no information about the parameter $\alpha$ is available, implying an effective mixture
$\bar{\rho}_{0}=\sum_{\alpha}p_{\alpha}\rho_{0}^{(\alpha)},$ (37)
and should also be compared to the case in which perfect knowledge of the initial distribution $\rho_{0}^{(\alpha)}$ is assumed.
After performing the noisy measurement of $\alpha$, the controller can gain knowledge of the initial distribution of the system. The conditional probability of obtaining a possible $\alpha$ value given the measurement result $m$ is ${p}(\alpha|m)=p_{\alpha}{p}(m|\alpha)/p(n)$, from which the updated system distribution ${\rho}_{X|m}(x)$ can be computed as:
@@
We now consider the combination of the situations of random initial distributions as described above and control limitations. Since each measurement outcome $m$ is unambiguously related to an optimal post-quench equilibrium distribution $\rho_{X|m}(\mathbf{x})$, our previous logic will apply. If there are $|\mathds{M}|$ possible measurements $m\in\{1,...,|\mathds{M}|\}$, and $N$ possible quenching Hamiltonians, $n\in\{1,...,N\}$, where $N<|\mathds{M}|$, then as before we need to design a repertoire and lookup table $n(m)$ mapping from the measurement results to the associated quenches in $\mathcal{H}$. Given a partition ${L}=\{\ell(n)\}_{n=1}^{N}$, similar calculations to those of Appendix A apply to obtain an optimal quenching Hamiltonian $H_{n}(\mathbf{x})=-k_{B}T\ln\rho_{n}^{\rm opt}(\mathbf{x})$ with:
$\rho_{n}^{\mathrm{opt}}(\mathbf{x})=\sum_{\alpha}p(\alpha|n)\rho^{(\alpha)}_{0}(\mathbf{x}),$ (40)
with $p(\alpha|n)=p_{\alpha}p(n|\alpha)/p(n)$, where $p(n)=\sum_{\alpha}p(n|\alpha)$, and $p(n|\alpha)=\sum_{m\in\ell(n)}P(m|\alpha)$. Details are reported in Appendix D, where we also show the equivalence of Eq. (40) above and Eq. (26), with use of the measurement-based expression Eq. (38). That is, we show that we may equivalently express the form of the optimal post-quench equilibrium $\rho^{\mathrm{opt}}_{n}$ as a probabilistic mixture of $\rho_{X|\mathbf{m}}$ across $\mathbf{m}$ (given $n$) or as a probabilistic mixture of $\rho_{0}^{(\alpha)}$ across $\alpha$ (given $n$).
Then likewise the average entropy production under control limitations reads:
$\langle S_{\mathrm{tot}}\rangle=\sum_{\alpha}p_{\alpha}\sum_{m}p(m|\alpha)D(\rho_{0}^{(\alpha)}||\rho_{n^{*}(m)}^{\rm opt}),$ (41)
which also exhibits a representation in terms of JSD and conditional mutual information akin to Eqs. (28) and (IV) respectively.
In particular, under the optimal action, $\langle S_{\mathrm{tot}}\rangle$ can be expressed as the average of a weighted multi-distributional Jensen-Shannon divergence, namely
$\langle S_{\mathrm{tot}}\rangle=\sum_{n}p(n)J_{p_{\alpha|n}}(\{\rho_{0}^{(\alpha)}\}_{\alpha}).$ (42)
The JSD in Eq. (42) at any given $n$ is the dissipation associated with action $n$ in response to the noisy measurement of $\alpha$. Accordingly, optimal partitions will cluster nearby distributions in the probability simplex (see Fig. 3 for an illustrative example).
Additionally, akin to Eq. (IV), we can re-express $\langle S_{\mathrm{tot}}\rangle$ in terms of a conditional mutual information between the system micro and macro states for a given a partition, $\langle S_{\mathrm{tot}}\rangle=I_{X,A;L}$, with $X,A,L$ denoting the random variables whose values are denoted $\mathbf{x},\alpha,n$ (see Appendix D for details). We note that here the macrostates $\alpha$ play the role of the measurement results $\mathbf{m}$ in expression (IV). Therefore, we can conclude that minimum dissipation in this scenario requires a minimum amount of extra correlations between microstates and macrostates, other than those strictly needed to perform the choice of the optimal action.
We demonstrate this scenario with a numerical example, obtaining the optimal partitions for a set of possible initial nonequilibrium distributions indexed by $\alpha=1,...,10$. We denote these distributions as $\{\rho^{(\alpha)}_{0}\}_{\alpha=1}^{10}$, with each $\rho_{0}^{(\alpha)}$ being a $3$-outcome distribution (for visualizability in the unit simplex). Moreover, we consider equal probabilities for each of them, $p_{\alpha}=1/10$ for all $\alpha$. We then numerically scan across all distinct partitionings of the $10$ distributions (macrostates) into $N$ response actions, computing the dissipation under the optimal choice in Eq. (40).
Figure 3: (a) Optimal partitions and post-quench equilibria for $N=3,5,7,9$ actions and $10$ possible initial distributions. The distributions $\{\rho_{0}^{(\alpha)}\}_{\alpha=1}^{10}$ are depicted as small dots, colored according to their assigned partition; large dots of corresponding colors represent the optimal quenching choices $\rho_{n}^{\mathrm{opt}}$. Additionally, regions are shaded based on the minimization of the KL divergence between $\{\rho_{n}^{\mathrm{opt}}\}_{n=1}^{N}$ and the points therein, akin to a Voronoi diagram Onishi and Imai (1997); Nielsen (2020). (b) Expected dissipation under optimal partition as a function of the number of actions $N$ when the round-truth distributions $\{\rho_{0}^{(\alpha)}\}_{\alpha=1}^{10}$ are sprinkled randomly into the unit $3$-simplex. $\mathbb{E}\langle S_{\mathrm{tot}}\rangle$ and $\Delta\langle S_{\mathrm{tot}}\rangle$ denote the mean and standard deviation of $\langle S_{\mathrm{tot}}\rangle$ in Eq. (41).
In Fig. 3a we depict the optimal partition for different number of allowed actions (memory size limitations) $N=3,5,7,9$. The different colored areas visually represent which quenching Hamiltonian will be chosen for the different distributions $\rho_{0}^{(\alpha)}$ (represented as small dots). The optimal quenching distributions for each partition, Eq. (40), are represented by the larger dots. In Fig. 3b we visualize the dissipation resulting from the optimal action repertoire with a number of available actions $N<10$. We show the mean $\mathds{E}[\langle S_{\mathrm{tot}}\rangle]$ and standard deviation $\Delta[\langle S_{\mathrm{tot}}\rangle]$ of the entropy production in the feedback control process across numerous samples of the choice of initial distributions. As can be observed, when $N$ grows from $1$ to $9$, the dissipation decreases confirming the optimal choice for the partitions.
@@ VI Conclusions and Discussion
In addition, such limitations are abundant in social systems. For example, many modern hunter-gatherer groups maintain social norms that help facilitate collective behavior, including the deterrence of free-riders and the encouragement of demand sharing (i.e., sharing food according to need; Lewis et al. (2014)). Thus, when an individual hunter-gatherer finds food or is successful in a hunt, their energy extraction actions are limited: they might eat some of the food immediately, but must save the rest to share with their camp. Further, while most hunter-gatherer populations store or preserve food, the extent of this storage is typically limited Armit and Finlayson (1992); Hamilton et al. (2024a). In contrast, individuals living in farming societies rely on being able to freely choose among multiple actions. Namely, to consume food immediately, or store it for later use. The increasingly complex technology available to farming and urbanized societies Hamilton et al. (2024b) suggests that having access to a larger action repertoire may be crucial to explaining the massive increase in per-person energy use between the ancient and modern world Krausmann et al. (2016).
There are many rich ways to extend the analysis in this paper, to capture even more features of real world systems that try to extract free energy from their environment. One obvious extension is to consider the case where there is more than just one agent who is extracting the free energy. In this extension the different agents would run quench-then-relax protocols on distributions involving different physical variables in the environment. It might also be that each agent observes some features of importance to some other ones, and so would need to communicate their information to the other agents, incurring a thermodynamic cost. Another avenue of future research arises from the fact that in this paper considering only a single agent we do not model the thermodynamic cost of the physical system that computes what action to take based on a given measurement. In particular we would need to consider such issues in a system that is used more than once, so that some of the variables in the agent need to be reinitialized at the beginning of each run.
@@ Acknowledgements
H. H., D. H. W., A. J. S., and C. K. acknowledge support from the Santa Fe Institute. D. W. acknowledges support from US NSF Grant CCF-2221345. G.  M. acknowledges support from the “Ramón y Cajal” program (No. RYC2021-031121-I), the CoQuSy project (No. PID2022-140506NB-C21), and the María de Maeztu project (No. CEX2021-001164-M) funded by MCIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR.
@@ Appendix B The case of perfect observation
Appendix C Details on the particle in $\mathbb{R}$ example
Herein we show that in the setting explored in the example of Sec. V.1 (1D, perfect observation, contiguous partition, $\rho_{0}$ has full support on $\mathbb{R}$), the unique way to achieve optimal quenching is to set partition boundaries at the positions given in Eq. (36).
Recall that an optimal partition must be probabilistically equitable, so that the distribution over measurements is maximally entropic: $S(p(m))=\log|\mathbb{M}|=\log N$. That is, we want the total probability of landing in $\ell(n)$ to be $1/N$, for each $n\in\{1,...,N\}$.
By the fact that $\rho_{0}(x)>0$ for all $x\in\mathbb{R}$, we have $\frac{d}{dx}F_{0}(x)>0$ for all $x\in\mathbb{R}$. Therefore $F_{0}(x)$ is monotonically increasing and uniquely invertible; we denote $F_{0}^{-1}:[0,1]\rightarrow\mathbb{R}$ as the inverse of $F_{0}$. Recall, $x_{0}=-\infty$, as we argued that our quenching Hamiltonian repertoire $\mathcal{H}$ should (collectively) reproduce the support of $\rho_{0}$. The location of $x_{1}$ is thus uniquely determined by the condition that $\mathbb{P}(X<x_{1})=1/N$. That is,
[…]
$\Rightarrow x_{1}=$ $F_{0}^{-1}\left(\frac{1}{N}\right).$
Now by induction we can obtain $x_{2},...,x_{N}$. Suppose the inductive hypothesis that for some $n\in\{2,...,N\}$ we have $x_{n-1}=F_{0}^{-1}(\frac{n-1}{N})$. Then, by monotonicity of $F_{0}$, and the optimality requirement of $\mathbb{P}(X\in\ell(n))=1/N$, the uniquely determined position of $x_{n}$ will satisfy
[…]
By hypothesis, $x_{n-1}=F_{0}^{-1}(\frac{n-1}{N})$, and thus $F_{0}(x_{n-1})=\frac{n-1}{N}$. Therefore, our requirement on $x_{n}$ (Eq. 61) becomes
$F_{0}(x_{n})-\frac{n-1}{N}$ $=\frac{1}{N}$ (62)
$\Rightarrow F_{0}(x_{n})$ $=\frac{n}{N},$
from which by solving for $x_{n}$ we obtain Eq. 36. Since we showed earlier that $x_{1}=F_{0}^{-1}(1/N)$ (in Eq. 60), we have by induction that Eq. 36 holds for all $n\in\{2,...,N\}$. (Including the case of $n=N$ in which we recover the already assumed $x_{N}=\infty$ at $F_{0}^{-1}(1)$, since $F(\infty)=1$.) See Fig. 2 in the main text for an example of the optimal partition boundary locations in the case where $\rho_{0}(x)$ is a standard normal density.
We stress that the uniqueness of this solution is lost if (a) the full support requirement fails, in which case regions of $\rho_{0}(x)=0$ may be present, corresponding to regions of $F_{0}(x)=\mathrm{const}$ (in which boundaries may be freely moved, retaining a partition identical in function). Uniqueness is also lost if (b) the contiguity requirement is dropped. We conjecture that this degeneracy vanishes in generic cases of noisy measurement, in which case localized noise will favor partitions with smaller boundaries.
Table of notation
Symbol Definition
$\mathcal{H}$ Set of quenching Hamiltonians $\{H_{k}\}_{k=1}^{N}$ for feedback control
${L}$ Set of partitions $\{l(n)\}_{n=1}^{N}$ grouping the measurement outcomes
$\rho_{0}(\mathbb{x})$ Initial marginal distribution of the system states $\mathbb{x}$
$\rho_{H}^{\rm eq}(\mathbf{x})$ Thermal equilibrium state for the system states $\mathbb{x}$ with Hamiltonian $H(\mathbb{x})$
$\rho_{X|\mathbb{m}}(\mathbb{x})$ Conditional distribution of the system states $\mathbb{x}$ given measurement result $\mathbb{m}$
$\rho_{n}^{\rm opt}(\mathbb{x})$ Distributions leading to optimal quenching Hamiltonians $H_{n}=-k_{B}T\log\rho_{n}^{\rm opt}$
$p(\mathbb{m})$ Marginal distribution of the measurement outcomes $\mathbb{m}$
$p_{M|\mathbb{x}}(\mathbb{m})$ Conditional distribution of measurement outcomes $\mathbb{m}$ for given system state $\mathbb{x}$
$n^{\ast}(\mathbb{m})$ Optimal choice for the quenching Hamiltonian given a measurement result $\mathcal{m}$
$p(n)$ Marginal distribution for obtaining partition $l(n)$ in the measurement
$p_{{L}|\mathbb{x}}(n)$ Conditional probability distribution for obtaining partition $l(n)$ given that the system is in state $\mathbb{x}$
$\rho_{0}^{\rm cg}(n)$ Marginal distribution for obtaining partition $l(n)$ in the noise-free scenario
$p_{\alpha}$ Probability of component $\alpha$ in the uncertain initial distributions scenario
$\rho_{0}^{(\alpha)}(\mathbb{x})$ Component $\alpha$ of the initial marginal distribution in the uncertain initial distributions scenario
$\bar{\rho}_{0}(\mathbb{x})$ Effective mixture of marginal distributions in the uncertain initial distributions scenario
$S(\rho)$ Shannon entropy of distribution $\rho(\mathbb{x})$
$D(\rho||\sigma)$ Kullback-Leibler divergence (relative entropy) between distributions $\rho(\mathbb{x})$ and $\sigma(\mathbb{x})$
$J_{\mathbf{p}}(\mathbf{P})$ Generalized Jensen-Shannon divergence among a weighted collection of distributions $\mathbf{P}$ with weights $\mathbf{p}$
$I_{X;M}$ Mutual information between the system states and measurement results
$I_{X;M|{L}}$ Conditional mutual information between the system and measurement results given a set of partitions ${L}$
$F_{H}^{\rm eq}$ Equilibrium free energy of the system for Hamiltonian $H(\mathbb{x})$
$\mathcal{F}_{H}(\rho)$ Nonequilibrium free energy of state $\rho$ for Hamiltonian $H(\mathbb{x})$
$\Delta\mathcal{F}_{H}$ Available nonequilibrium free energy in the initial state, $\mathcal{F}_{H}(\rho_{0})-F_{H}^{\rm eq}$
$S_{\mathrm{tot}}(\mathbb{m})$ Entropy production (free energy loss) in the feedback control process for measurement outcome $\mathbb{m}$
$\langle S_{\mathrm{tot}}\rangle$ Average entropy production (average free energy loss) in the feedback control process after many runs
$W_{\rm ideal}(\mathbf{m})$ Work extracted in the ideal feedback control protocol for measurement outcome $\mathbb{m}$
$\langle W_{\rm ideal}\rangle$ Average work extracted in the ideal feedback control protocol after many runs
$W(\mathbf{m})$ Work extracted in the general feedback control protocol with limitations for measurement outcome $\mathbb{m}$
$\langle W\rangle$ Average work extracted in the general feedback control protocol after many runs
$W_{\rm cost}^{q}$ Work cost for performing a predefined quench $H_{q}(\mathbb{x})$
$\langle W_{\rm cost}\rangle$ Average work cost for performing the quenches after many runs
$\langle\Delta W\rangle$ Average work gain from feedback control with respect to a deterministic quench after many runs
$\langle\Delta W\rangle_{r}$ Average work gain from feedback control with respect to random quenching using distribution $r(q)$
Table 1: Table of notation for the main symbols used in the paper.
Appendix D Additional calculations for an arbitrary random initial distribution
Herein we provide details on the derivation of results for the scenario of feedback action based on a noisy measurement of a distribution-specifying parameter $\alpha$, considered in Section V.2. The parameter $\alpha$, sampled with probability $p_{\alpha}$, specifies the initial nonequilibrium distribution $\rho_{0}^{(\alpha)}$. Then a noisy measurement $m$ of $\alpha$ is produced with probability $p(m|\alpha)$, after which the feedback action $n=n^{\star}(m)$ is employed in response. Here again a lookup table $n^{\star}$ determines a partition $\{\ell(n)\}_{n}$ by $m\in\ell(n)\Rightarrow n^{\star}(m)=n$.
First we note that the total entropy production in Eq. (41) can be re-expressed as follows
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{\alpha}p_{\alpha}\sum_{m}P(m|\alpha)D(\rho^{(\alpha)}_{0}||\rho_{n(m)}^{\mathrm{opt}})$ (63)
$=\sum_{\alpha}p_{\alpha}\sum_{n}p(n|\alpha)D(\rho^{(\alpha)}_{0}||\rho_{n}),$
with $p(n|\alpha)=\sum_{m\in\ell(n)}P(m|\alpha)$. Let us then define the conditional probability that parameter $\alpha$ was sampled given that action $n$ is taken
$p_{\alpha|n}=\frac{p(n|\alpha)p_{\alpha}}{p(n)},$ (64)
with $p(n)=\sum_{\alpha}p_{\alpha}p(n|\alpha)$ the marginal probability of taking action $n$. Then continuing from Eq. (63) we obtain
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{\alpha,n}p_{\alpha}p(n|\alpha)D(\rho^{(\alpha)}_{0}||\rho_{n}^{\mathrm{opt}})$ (65)
$=\sum_{n}p(n)\sum_{\alpha}p_{\alpha|n}D(\rho^{(\alpha)}_{0}||\rho_{n}^{\mathrm{opt}}).$
The optimal post-quench equilibrium $\rho_{n}^{\mathrm{opt}}$ minimizes the summand $\sum_{\alpha}p_{\alpha|n}D(\rho^{(\alpha)}_{0}||\rho_{n}^{\mathrm{opt}})$ in the above Eq. (65), and thus by the derivation of Appendix A we have that the solution has a probabilistic mixture form:
$\rho_{n}^{\mathrm{opt}}=\sum_{\alpha}p_{\alpha|n}\rho^{(\alpha)}_{0}.$ (66)
which was reported as Eq. (40) in the main text. Note the difference between Eq. (26) and Eq. (66) above; we will reconcile these below [see Eq. (79)].
In the optimal scenario of Eq. (40), the dissipated work becomes
$\langle S_{\mathrm{tot}}\rangle$ $=\sum_{n}p(n)\sum_{\alpha}p_{\alpha|n}D\left(\rho^{(\alpha)}_{0}\left|\left|\sum_{\alpha}p_{\alpha|n}\rho^{(\alpha)}_{0}\right.\right.\right)$ (67)
$=\sum_{n}p(n)\sum_{\alpha}p_{\alpha|n}\sum_{x}\rho_{0}^{(\alpha)}(\mathbf{x})\log\frac{\rho_{0}^{(\alpha)}(\mathbf{x})}{\rho_{n}^{\mathrm{opt}}(\mathbf{x})}$
$=\sum_{n}p(n)\left(S(\rho_{n}^{\mathrm{opt}})-\sum_{\alpha}p_{\alpha|n}S(\rho_{0}^{(\alpha)})\right)$
$=\sum_{n}p(n)J_{p_{\alpha|n}}(\{\rho_{0}^{(\alpha)}\}_{\alpha}).$
The latter expresses that the dissipation is the average weighted multi-distributional JSD among the possible ground truth distributions $\{\rho_{0}^{(\alpha)}\}_{\alpha}$ as weighted by the posterior probability of $\alpha$ given $n$.
Additionally, we can re-express $\langle S_{\mathrm{tot}}\rangle$ in terms of a conditional mutual information. The conditional mutual information among three random variables $X,Y,Z$ is given by the general expression
$I_{X,Y;Z}=\sum_{z}P(z)\sum_{x,y}P(x,y|z)\log\frac{P(x,y|z)}{P(x|z)P(y|z)}.$ (68)
We will show below that $\langle S_{\mathrm{tot}}\rangle=I_{X,A;\mathcal{L}}$, with shorthand $X,A,\mathcal{L}$ for the random variables whose values are being denoted $\mathbf{x},\alpha,n$. First we define the joint distribution of $(\alpha,\mathbf{x})$ given $n$, denoted
$p(\alpha,\mathbf{x}|n)$ $=p_{\alpha|n}\rho_{0}^{(\alpha)}(\mathbf{x}).$ (69)
The conditional distribution of $\mathbf{x}$ given $n$ is
$\sum_{\alpha}p(\alpha,\mathbf{x}|n)$ $=\sum_{\alpha}p_{\alpha|n}\rho_{0}^{(\alpha)}(\mathbf{x})=\rho_{n}^{\mathrm{opt}}(\mathbf{x}).$ (70)
The overall probability of action $n$ is
$p(n)$ $=\sum_{\alpha,m}p_{\alpha}P(m|\alpha)\delta_{n,n^{\star}(m)}$ (71)
$=\sum_{\alpha}p_{\alpha}\sum_{m\in\ell(n)}P(m|\alpha)$
$=\sum_{\alpha}p_{\alpha}p(n|m),$
that is, it equals the optimal distribution for the quenching Hamiltonian in Eq. (66). Therefore, we have
$I_{X,A;L}$ $=\sum_{n}p(n)\sum_{\alpha,\mathbf{x}}p(\alpha,\mathbf{x}|n)\log\frac{p(\alpha,\mathbf{x}|n)}{p_{\alpha|n}\rho_{n}^{\mathrm{opt}}(\mathbf{x})}$ (72)
$=\sum_{n}p(n)\left(\sum_{\alpha,\mathbf{x}}p_{\alpha|n}\rho_{0}^{(\alpha)}(\mathbf{x})\log\frac{\rho_{0}^{(\alpha)}(\mathbf{x})}{\sum_{\alpha}p_{\alpha|n}\rho_{0}^{(\alpha)}(\mathbf{x})}\right)$
$=\sum_{n}p(n)\left(S(\rho_{n}^{\mathrm{opt}})-\sum_{\alpha}p_{\alpha}S(\rho_{0}^{(\alpha)})\right),$
reproducing Eq. (67). Thus the mean dissipated work is the conditional mutual information between the post-quench equilibrium state $\mathbf{x}$ and the ground truth index $\alpha$, with conditionality on the action taken $n$.
As a consistency check, let’s reconcile Eq. (40) with Eq. (26) of the manuscript, to demonstrate the compatibility of the measurement-based expressions with the case of arbitrary initial distributions. Given the measurement of $m$, we have a posterior over $\alpha$ of the form
$p(\alpha|m)$ $=\frac{P(m|\alpha)p_{\alpha}}{p(m)}$ (73)
$=\frac{P(m|\alpha)p_{\alpha}}{\sum_{\alpha}P(m|\alpha)p_{\alpha}},$
from which we can calculate the post-measurement distribution as
Thus by following the reasoning of Appendix A, the $n$th optimal action should result in a post-quench equilibrium of the form:
$\rho_{n}^{\mathrm{opt}}(\mathbf{x})$ $=\sum_{m}p(m|n)\rho_{X|m}(\mathbf{x}).$ (75)
Eq. (75) is of the form of a mixture distribution across measurement outcomes, as opposed to the expression derived above, Eq. (66), which is a mixture distribution across the hidden parameter $\alpha$. We next show the equivalence of the two.
Rewriting Eq. (75),
$\rho_{n}^{\mathrm{opt}}(\mathbf{x})$ $=\sum_{m}p(m|n)\rho_{X|m}(\mathbf{x})$ (76)
$=\sum_{m}\frac{\mathds{1}\{m\in\ell(n)\}p(m)}{p(n)}\rho_{X|m}(\mathbf{x})$
$=\frac{\sum_{m\in\ell(n)}p(m)\rho_{X|m}(\mathbf{x})}{\sum_{m\in\ell(n)}p(m)}.$
Now noting that $p(m)=\sum_{\alpha}P(m|\alpha)p_{\alpha}$, we have
$\sum_{m\in\ell(n)}p(m)$ $=\sum_{m\in\ell(n)}\sum_{\alpha}P(m|\alpha)p_{\alpha}$ (77)
$=\sum_{\alpha}\left(\sum_{m\in\ell(n)}P(m|\alpha)\right)p_{\alpha}$
$=\sum_{\alpha}p(n|\alpha)p_{\alpha}=p(n).$
Similarly, we also have
$\sum_{m\in\ell(n)}p(m)\rho_{X|m}(\mathbf{x})$ $=\sum_{m\in\ell(n)}p(m)\left(\sum_{\alpha}p(\alpha|m)\rho_{0}^{(\alpha)}(\mathbf{x})\right)$ (78)
$=\sum_{m\in\ell(n)}\sum_{\alpha}P(m|\alpha)p_{\alpha}\rho_{0}^{(\alpha)}(\mathbf{x})$
$=\sum_{\alpha}p_{\alpha}\rho_{0}^{(\alpha)}(\mathbf{x})\sum_{m\in\ell(n)}P(m|\alpha)$
$=\sum_{\alpha}\rho_{0}^{(\alpha)}(\mathbf{x})p(n|\alpha)p_{\alpha},$
Therefore, replacing the above two expressions on Eq. (76) we obtain:
$\rho_{n}^{\mathrm{opt}}(\mathbf{x})$ $=\sum_{m}p(m|n)\rho_{X|m}(\mathbf{x})$ (79)
$=\frac{\sum_{\alpha}\rho_{0}^{(\alpha)}(\mathbf{x})p(n|\alpha)p_{\alpha}}{p(n)}$
$=\sum_{\alpha}\rho_{0}^{(\alpha)}(\mathbf{x})p(\alpha|n),$
with $p(\alpha|n)=p(n|\alpha)p_{\alpha}/p(n)$.
D.1 Relation to original measurement scheme
Finally, we show how the original measurement scenario considered in Sec. II is equivalent to the random initial distributions scenario when $\alpha\sim p_{\alpha}$ is assumed to be a random microstate $\mathbf{x}\sim\rho_{0}$, again passed through a noisy channel [$P(\mathbf{m}|\mathbf{x})$ rather than $P(m|\alpha)$], and with the associated set of ground truth distributions $\{\rho_{0}^{(\alpha)}\}_{\alpha}$ being set equal to $\{\delta_{\mathbf{x},\mathbf{x}^{\prime}}\}_{\mathbf{x}}$.
Under these circumstances ($\alpha\rightarrow\mathbf{x},p_{\alpha}\rightarrow\rho_{0}(\mathbf{x}),\rho_{0}^{(\alpha)}(\mathbf{x})\rightarrow\rho_{0}^{(\mathbf{x})}(\mathbf{x}^{\prime})=\delta_{\mathbf{x},\mathbf{x}^{\prime}}$), we have
$p(\alpha|m)\rightarrow p(\mathbf{x}|\mathbf{m})$ $=\frac{P(\mathbf{m}|\mathbf{x})\rho_{0}(\mathbf{x})}{\sum_{\mathbf{x}}P(\mathbf{m}|\mathbf{x})\rho_{0}(\mathbf{x})}$ (80)
$=\frac{P(\mathbf{m}|\mathbf{x})\rho_{0}(\mathbf{x})}{p(\mathbf{m})},$
and thus the posterior is
$\rho_{X|\mathbf{m}}(\mathbf{x}^{\prime})$ $=\sum_{\alpha}p(\alpha|\mathbf{m})\rho_{0}^{(\alpha)}(\mathbf{x}^{\prime})$ (81)
$=\sum_{\mathbf{x}}\frac{P(\mathbf{m}|\mathbf{x})\rho_{0}(\mathbf{x})}{p(\mathbf{m})}\delta_{\mathbf{x},\mathbf{x}^{\prime}}$
$=\frac{P(\mathbf{m}|\mathbf{x})\rho_{0}(\mathbf{x}^{\prime})}{p(\mathbf{m})}$
i.e., the posterior with respect to the initial microstate distribution $\rho_{0}$. Note that in this scenario, the initial average microstate distribution, given in general by Eq. 37, becomes identical to the microstate distribution itself:
$\bar{\rho}_{0}=\sum_{\mathbf{x}_{0}}\rho_{0}(\mathbf{x}_{0})\delta_{\mathbf{x}_{0},\mathbf{x}}=\rho_{0}.$ (82)